(34) Continued... Suppose X and Y have geometric distributions with different parameters Geometric(\alpha) and Geometric(\beta). Find the convolution of their marginal distribution1.

1. n-11-a





n-11-a
Answer #1

2-1 r=1(by law of total probability)

= \sum_{x=1}^{n-1}P(x+Y=n)P(X=x)

2-1 r=1

2-1 r=1 (since distribution of X and Y are geometric, given)

= \alpha \beta (1-\beta )^{n-2}\sum_{x=1}^{n-1}\frac{(1-\alpha)^{x-1}}{(1-\beta )^{x-1}}

72-1 1-B 1-5

α-β

where , n=2,3,.......